# Product Design Using a Generative Adversarial Network: Incorporating Consumer Preferences and External Data

### v2

@@ 2.1 Overview of the Conceptual Model
A key contribution of our proposed framework is its introduction of a conditional GAN to marketing applications. One challenge is that conditional GAN models typically use categorical labels as additional inputs, while the predicted popularity measures in our context are continuous class labels. To address this issue, we employ the CcGAN that was introduced by Ding et al., (2020). The CcGAN is an extension of the conditional GAN model, and it was designed to generate images based on continuous, scalar conditions. By redefining the conventional conditional GAN approach to make it suitable for continuous scenarios and introducing a novel method that can integrate continuous labels into the generation process, the CcGAN outperforms the conditional GAN in terms of effectively handling continuous labels.
@@
Another limitation of the enhanced model is that it does not fully utilize the information contained in external UGC. By nature, UGC data inherently contain rich consumer preference information, as users voluntarily generate the content. In our context, i.e., photo templates, the UGC on social media and image websites usually contains a vast set of photos taken by individual users in front of various backgrounds. This type of image represents not only a rich trove of product design data (i.e., the backgrounds in the case of photo templates) but also a valuable source of consumer preference information. The fact that an individual user chose to take photos in front of a particular background means that the user likes the background. As such, the external images can also be used to train the predictor model to understand how consumers evaluate new features in external product designs.
@@
The advanced model follows a workflow similar to that of the enhanced model except for the fact that the predictor model is trained on both internal and external consumer preferences. Following the same logic as that used for the internal data, the predictor model learns the mappings between the users’ appearances and the backgrounds in the UGC images. Upon pooling the mappings between the consumer appearances and product designs in both the internal and external datasets, the trained prediction model captures a larger set of consumers and their preferences over a wider variety of product designs. The predictor model is labeled Prediction Model 2 in Figure 1. Compared with Prediction Model 1, this version can recognize popular out-of-sample features from external designs.
@@
Comparing the three models illustrates the benefits of incorporating external designs and consumer preference information into the product design generation process. The full version of our proposed model (the CcGAN with internal and external consumer preferences) is a semi-supervised deep generative framework that creates consumer-preferred product designs in an automatic and cost-effective manner. It has several highlights. First, different from previous studies in which consumer preference information was only used to train predictor models, we incorporate consumer preference information into both the predictor model and the generator model. In particular, the popularity predicted for each product serves as a label and an additional input for the generator, which allows the generator to create products that are more likely to be preferred by the target consumers. Second, we exploit the UGC derived from external data sources. The various product designs contained in UGC enrich the generative model and allow it to incorporate new product features that are not included in the internal data. The rich consumer preference information embedded in the UGC improves the supervision of the generative model and allows it to recognize popular new product features that are not presented to the internal consumers.
We contribute to the literature concerning deep learning models in marketing. Most studies in this area use unstructured data to generate structured firm insights. For instance, Liu et al., (2020) collected unstructured image data from social media and used a deep convolutional neural network (CNN) to extract and predict consumer perceptions of brand attributes. Our work differs in that we employ unstructured data to directly generate unstructured firm decisions (i.e., product designs as images) using state-of-the-art deep learning models. Our proposed framework goes beyond the descriptive task and provides prescriptive solutions for the product design decisions of firms.
We also contribute to the nascent literature on deep generative models in marketing. A few works have used deep generative models to create new product designs. For example, Dew et al., 2022 used multiview representation learning to design brand logos. Burnap et al., 2023 proposed a model to augment the commonly used aesthetic design process by predicting aesthetic scores and automatically generating innovative and appealing product designs; the predictor model and the generator models were separately trained. Our work differs in that we incorporate consumer preference information into the generation process so that the predictor model guides the generative model. We also introduce the CcGAN model to the marketing field, allowing us to systematically incorporate consumer preferences in an automated and large-scale manner. In addition, while previous works generated one aspect of product designs such as logos (Dew et al.,, 2022) or the aesthetic appearances of cars (Burnap et al.,, 2023), we directly design the entire product (i.e., the photo templates are the products of the firm). This setting allows us to abstract away from other design factors that might influence consumer choices.
2.2 Prediction Model
@@
This subsection describes the detailed procedure for training the prediction model, as shown in Equation 2.1. The primary inputs of the prediction model consist of consumer appearances (i.e., consumer faces in our empirical context) and product designs (i.e., photo templates or backgrounds in our empirical context). Given the high dimensionality of the input data, we preprocess the given images by embedding them into a lower-dimensional vector space. For consumer faces, we follow the deep neural network implemented by the OpenFace project (Amos et al.,, 2016). As illustrated in Figure A.1 in the appendix, this architecture was trained for facial recognition. It first detects and crops faces from a given image. Then, it produces a 128-dimensional intermediate layer through a deep neutral network, which represents a low-dimensional embedding of the face image. For the product design images, we adopt the Caffe BVLC reference model (Jia et al.,, 2014) to obtain image embeddings. This model uses the CNN designed by Alex Krizhevsky (AlexNet) with some modifications, as shown in Figure A.2 in the appendix. It provides pretrained general networks that can be used to obtain low-dimensional embeddings of image characteristics. Using the pretrained weights of the Caffe BVLC reference model, we can directly go to fully connected (fc) layer 8 to obtain a 1000-dimensional intermediate layer that captures the visual characteristics of the input product design image.
@@
$P_{j}(\widehat{\theta})=\frac{1}{N}\sum_{i\in\mathcal{I}}p_{ij}(\widehat{\theta}),$ (2.2)
where $\mathcal{I}$ denotes the total consumer set and $N$ is the size of the consumer base. The aggregate popularity measure is later incorporated into the constructed generative models to guide the process of generating new product designs.
As discussed in Section 2.1, we train two prediction models: one with internal preferences only and another with both internal and external preferences. We train these two prediction models via similar training processes. The difference is that Prediction Model 1 is trained with internal consumer appearances and product designs only, while Prediction Model 2 is trained using both internal and external consumer appearances and product designs. Note that external consumers are used in the training stage of Prediction Model 2 but not in the prediction stage; when using the trained models to calculate the aggregate popularity measures, we average over the choice probabilities of the internal consumers only for both Prediction Models 1 and 2. This is because we are interested in predicting whether the internal consumers of firms prefer the new template designs. The external consumer preference information still enhances the performance of Prediction Model 2 by training the model to recognize popular design features in external designs. The underlying assumption is that external consumers with similar appearances to those of the internal consumers share similar preferences. Therefore, learning the preferences of the external consumers helps us infer the preferences of the internal consumers for the new design features acquired from the external designs.
2.3 Generative Models
@@
When training the baseline model without consumer preferences, we adopt the standard and well-established model architecture of the DCGAN method (Radford et al.,,2015) to generate new images. This model takes the internal and external product images as real images input and identifies a set of parameters that enables the samples generated from random noises to closely resemble the training data. As the training process is unsupervised and lacks consumer preference information, the generated images are not necessarily preferred by consumers.
When training the enhanced model with internal consumer preferences, we estimate a conditional GAN model and let the internal consumer preferences guide the generator to produce appealing designs that are tailored to the targeted consumers. Specifically, we use the CcGAN developed in Ding et al., (2020) to handle continuous labels.
@@
$y_{j}=P_{j}(\widehat{\theta}_{1}),$ (2.3)
where $\widehat{\theta}_{1}$ is the estimated parameter of Prediction Model 1 and $P_{j}(\widehat{\theta}_{1})$ is defined in Equation 2.2. This popularity measure serves as a continuous class label and is fed into both the generator and the discriminator in the CcGAN to guide the generator when creating new designs. Specifically, following Ding et al., (2020), we define the hard vicinal discriminator loss (HVDL) as
^$\begin{split}\widehat{\mathcal{L}}(D)=&-\frac{C_{1}}{M^{r}}\sum_{j=1}^{M^{r}}\sum_{i=1}^{M^{r}}\mathbb{E}_{\epsilon^{r}\sim N(0,\sigma^{2})}\Big{[}\frac{\mathbbm{1}_{\{|y_{j}^{r}+\epsilon^{r}-y_{i}^{r}|\leq\kappa\}}}{N_{y_{j}^{r}+\epsilon^{r},\kappa}^{r}}\mbox{log}(D(\boldsymbol{x}_{i}^{r},y_{j}^{r}+\epsilon^{r}))\Big{]}\\ &-\frac{C_{2}}{M^{g}}\sum_{j=1}^{M^{g}}\sum_{i=1}^{M^{g}}\mathbb{E}_{\epsilon^{g}\sim N(0,\sigma^{2})}\Big{[}\frac{\mathbbm{1}_{\{|y_{j}^{g}+\epsilon^{g}-y_{i}^{g}|\leq\kappa\}}}{N_{y_{j}^{g}+\epsilon^{g},\kappa}^{r}}\mbox{log}(1-D(\boldsymbol{x}_{i}^{g},y_{j}^{g}+\epsilon^{g}))\Big{]},\end{split}$ (2.4)
where $C_{1}$ and $C_{2}$ are constants, $\boldsymbol{x}_{i}^{r}$ and $\boldsymbol{x}_{i}^{g}$ stand for real and generated images, respectively, $y_{i}^{r}$ and $y_{i}^{g}$ are the popularity labels of $\boldsymbol{x}_{i}^{r}$ and $\boldsymbol{x}_{i}^{g}$, and $\epsilon^{r}=y-y_{j}^{r}$ and $\epsilon^{g}=y-y_{j}^{g}$. $N_{y_{j}^{r}+\epsilon^{r},k}^{r}$ is the number of the $y_{i}^{r}$ that satisfies $|y_{j}^{r}+\epsilon^{r}-y_{i}^{r}|\leq\kappa$. The loss function for the generator is defined as
^$\widehat{\mathcal{L}}(G)=-\frac{1}{M^{g}}\sum_{i=1}^{M^{g}}\mathbb{E}_{\epsilon^{g}\sim N(0,\sigma^{2})}\mbox{log}(D(G(\boldsymbol{z}_{i},y_{i}^{g}+\epsilon^{g}),y_{i}^{g}+\epsilon^{g})),$ (2.5)
@@
In summary, combining Equations 2.3, 2.4, and 2.5, we integrate consumer preferences into the generation process by directly incorporating the popularity measure $P_{j}(\widehat{\theta})$ from the predictor model into the loss functions of both the discriminator and the generator of the CcGAN. This setup allows us to instruct the generator to create new images with a specific popularity level.
The training process of advanced model with both internal and external consumer preferences follows a procedure similar to that used for training the enhanced model. The only difference is that we use the aggregate popularity measures from Prediction Model 2, $P_{j}(\widehat{\theta}_{2})$, as additional input labels to be fed into the generator and the discriminator. As Prediction Model 2 is trained on both internal and external consumer choices, the new CcGAN is able to incorporate new design features from the external sources that are appealing to the internal consumers.
3 Institutional Setting and Data
3.1 Institutional Background
As an empirical application, we apply our proposed framework to a real business setting: a large self-executed photography chain in China. The company operates self-executed photo booths in approximately twenty shopping malls in major cities. The photo booths allow consumers to choose photo templates, take photos with the chosen templates, and print their photos. The photo booths are self-operated, so consumers follow the instructions on the photo machines placed in the booth. Figure 2 shows the interface of such a photo machine. Four photo categories can be selected: main (single), friends, family, and lovers. We focus on the “main” category in our analysis, which is designed for one consumer to take photos.
@@ 3.1 Institutional Background
3.2 Internal Data
The internal dataset obtained from the company contains two parts. The first part includes the original template images that the company offered to customers, without the presence of consumers inside the images. The company offered 83 photo themes in total during the observation period. Each theme contains 1 to 12 templates, yielding a total of 585 templates across all the themes. The second part of the internal dataset contains the complete order histories of 2195 consumers who had taken at least one photo between October 2017 and August 2018. For each consumer, we observe their selected themes and templates and the photos taken with the chosen templates. We also observe the ranking orders of the themes offered on the screen at the times they made their choices.
The internal data patterns suggest that the templates vary significantly in popularity, potentially due to different design features and heterogeneous consumer preferences. Figure 4 presents the number of consumers who choose each template in the graffiti theme and the traditional Chinese theme. The X-axis is the template index, which represents the order in which the templates belonging to each theme are displayed on the screen. As shown in the figure, the templates have very different popularity levels both across themes and within the same theme. For graffiti, the number of adoptions for the most-popular template is four times that of the least-popular template. For the traditional Chinese theme, the least-preferred template has zero adoptions, whereas the most-preferred template has more than 60 adoptions. The variation in template popularity, especially the presence of low-popularity templates, suggests that the company needs to improve the attractiveness of its template designs. As another piece of evidence, the internal dataset shows that only 2195 out of the 6038 registered consumers ended up taking photos in the booth after browsing the photo template options. Among the 2195 consumers, fewer than 100 consumers made multiple store visits, and only 15% of the consumers had repeated time purchases. These numbers suggest that the company can benefit from incorporating consumer preferences to generate better designs that are appealing to consumers.
Figure 4: Number of Adopters of Each Template Under Different Themes
Refer to caption Refer to caption
(a) Graffiti Theme (b) Traditional Chinese Theme
3.3 External Data
@@
After gathering the external images with human objects and backgrounds, we preprocess the images before they are used as inputs for the proposed model. Specifically, training the model requires a set of pure templates without human objects and a set of consumer face images, while the raw external images we collect are combinations of template backgrounds and consumer faces, so we need to separate the two components. For each external image, we remove the human object in it and inpaint the remaining part of the image to obtain a complete background image, which is similar to the pure template images contained in the internal dataset. Additionally, we save the cropped human faces as new images, which are similar to the consumer face images in the internal dataset. Figure 5 presents examples of the internal templates, raw external photos, inpainted template images, and cropped consumer face images, all of which are from the graffiti theme. As shown in the graphs, the collected external images are similar in style to the internal images and have sufficient variations. The inpainting process works well and can generate high-quality templates and consumer face images.
@@
4 Applying the Model to Practice
@@
We apply the proposed model to generate new product designs for the unmanned photo gallery firm. This section describes the implementation details, and the next section presents the results.
4.1 Training the Prediction Model
@@
As discussed in Section 2, the feature embeddings of consumer faces and background templates serve as the inputs for the prediction model. Regarding the consumer face features, most consumers took more than one photo, so multiple face images are available for each consumer. We first obtain the embedding of each face image. We then average across the embeddings of the same consumer to obtain a 128-dimensional facial vector $\overrightarrow{X}_{i}$ for that consumer. With respect to the template features, in addition to the 1000-dimensional embeddings of the visual characteristics $\overrightarrow{V}_{j}$, we also include the display ranking of each template on the screen $\overrightarrow{Z}_{j}$, as its display ranking affects its probability of being chosen. Concatenating $\overrightarrow{X}_{i}$, $\overrightarrow{V}_{j}$, and $\overrightarrow{Z}_{j}$ yields a 1129-dimensional feature input for training the prediction model.
@@
The output variable of the prediction model is constructed from the consumers’ revealed preferences, namely, their template choices. In our setting, consumers first choose a theme among all themes; only one template from each theme is displayed as the cover picture for each theme at this stage. Once consumers click on the chosen theme, they can browse all templates in the chosen theme and choose some of the templates to take photos. We assume that consumers make binary choices for each template they browse and can choose multiple templates since consumers can select an unlimited number of themes and templates within the paid time window. The chosen templates belonging to the chosen theme naturally serve as the positive samples. The unchosen templates belonging to the chosen theme serve as the negative samples because consumers see them but decide not to take photos with them. For the unchosen themes, the cover-page templates count toward the negative samples because consumers see them upon choosing the themes and do not choose them. However, the other templates belonging to the unchosen themes do not count as negative samples and are not included in the analysis because consumers do not have a chance to see them. Overall, the output variable $Y_{ij}$ for consumer $i$ and template $j$ is constructed as follows:
@@
$Y_{ij}=\left\{\begin{array}[]{ll}1&\text{if template }j\text{ is chosen by consumer }i\\ 0&\text{if template }j\text{ is not chosen but belongs to a chosen theme}\\ 0&\text{if template }j\text{ is the cover-page template of an unchosen theme }\end{array}\right..$ (4.1)
@@
Given the definitions of the input and output variables, we estimate Prediction Model 1 using internal consumer preferences only. We focus on the photo samples from the category “Main (single)”, which contains 931 consumers and 33,804 consumer-template pairs. Each consumer-template pair corresponds to one consumer choice. Among these 33,804 choices, 3,629 are positive samples (i.e., consumers choose to take photos with the templates), and 30,175 are negative samples. We retain $40\%$ of the observations as the test set and use the remaining data for training the model. The training and test sets include 20,282 and 13,522 observations, respectively. As positive samples only account for 10.7% of the total sample size, an unbalanced sample problem is encountered. To alleviate this problem, we upsample the minority class to 18,000 to better match the majority class. We concatenate the upsampled minority class and the majority class to form the final training set and train the model using a random forest. The number of trees in the forest is set to 100. The class weights are set to be balanced.
@@
We estimate Prediction Model 2 using both internal and external consumer preferences. Specifically, external UGC data contain photos taken by individual users in front of various backgrounds. Each photo represents a user’s chosen background and thus serves as a positive sample for the prediction model. There are 1375 external photos or choice observations. Combining these with the 33,804 internal observations yields a sample size of 35,179 photos. We use 70% of the observations, or 24,625 photos, as the training set. Among them, 2,544 are positive samples, and 22,081 are negative samples. To address the unbalanced sample issue, we upsample the minority class to 20,000 to align with the majority class. We estimate Prediction Model 2 using a random forest again, maintaining the same hyperparameters as those employed in Prediction Model 1.
4.2 Training the Generative Models and Generating New Designs

### v3

@@ Introduction
A key contribution of our proposed framework is that we introduce the conditional GAN to marketing applications and adapt it to suit our context. First, conditional GAN models typically use categorical labels as additional inputs, while the predicted popularity measures in our context are continuous class labels. To address this issue, we employ the CcGAN that was introduced by Ding et al., (2020). The CcGAN is an extension of the conditional GAN model, and it was designed to generate images based on continuous, scalar conditions. By redefining the conventional conditional GAN approach to make it suitable for continuous scenarios and introducing a novel method that can integrate continuous labels into the generation process, the CcGAN outperforms the conditional GAN in terms of effectively handling continuous labels.
Second, rather than using CcGAN from Ding et al., (2020) as-is, we modify it to suit our context by accounting for the ordinal nature of popularity labels. In Ding’s applications, the continuous labels represent yaw angles of 3-D chair models and human ages, which do not imply preference (e.g., age 15 is neither better nor worse than age 30). However, in our case, higher popularity levels indicate greater consumer preference; we aim to generate designs with higher popularity. To address this, we adapt CcGAN to recognize the ordinal nature of popularity labels by modifying its loss function to include label weights. Further details are provided in the following subsections.
@@
Another limitation of the enhanced model is that it does not fully utilize the information contained in external UGC. By nature, UGC data inherently contain rich consumer preference information, as users voluntarily generate the content. In our context of photo templates, the UGC on social media and image websites usually contains a vast set of photos taken by individual users in front of various backgrounds. This type of image represents not only a rich trove of product design data (i.e., the backgrounds in the case of photo templates) but also a valuable source of consumer preference information. The fact that an individual user chose to take photos in front of a particular background means that the user likes the background. As such, the external images can also be used to train the predictor model to understand how consumers evaluate new features in external product designs.
@@
The advanced model follows a workflow similar to that of the enhanced model except for the fact that the predictor model is trained on both internal and external consumer preferences. Following the same logic as that used for the internal data, the predictor model learns the mappings between the users’ appearances and the backgrounds in the UGC images. Upon pooling the mappings between the consumer appearances and product designs in both the internal and external datasets, the trained prediction model captures a larger set of consumers and their preferences over a wider variety of product designs. The predictor model is labeled “Prediction Model 2” in Figure 1. Compared with Prediction Model 1, this version can recognize popular out-of-sample features from external designs.
@@
Comparing the three models highlights the advantages of integrating external designs and consumer preferences into product design generation. Our full model, CcGAN with internal and external consumer preferences, is a semi-supervised deep generative framework that automatically and cost-effectively creates consumer-preferred designs. Unlike previous studies that used consumer preference data solely for predictor models, we incorporate this data into both predictor and generator models. Predicted popularity serves as both a label and an input for the generator, enabling it to create products likely to appeal to target consumers. Additionally, we utilize UGC from external sources to enrich the generative model with diverse designs and new features absent from internal data. The extensive consumer preference information in UGC enhances the model’s ability to identify popular features not evident in internal consumer data.
Prediction Model
@@
This subsection describes the detailed procedure for training the prediction model, as shown in Equation 1. The primary inputs of the prediction model consist of consumer appearances (i.e., consumer faces in our empirical context) and product designs (i.e., photo templates or backgrounds in our empirical context). Given the high dimensionality of the input data, we preprocess the given images by embedding them into a lower-dimensional vector space. For consumer faces, we follow the deep neural network implemented by the OpenFace project (Amos et al.,, 2016). As illustrated in the online appendix, this architecture was trained for facial recognition. It first detects and crops faces from a given image. Then, it produces a 128-dimensional intermediate layer through a deep neutral network, which represents a low-dimensional embedding of the face image. For the product design images, we adopt the Caffe BVLC reference model (Jia et al.,, 2014) to obtain image embeddings. This model uses the CNN designed by Alex Krizhevsky (AlexNet) with some modifications, as shown in the online appendix. It provides pretrained general networks that can be used to obtain low-dimensional embeddings of image characteristics. Using the pretrained weights of the Caffe BVLC reference model, we can directly go to fully connected (fc) layer 8 to obtain a 1000-dimensional intermediate layer that captures the visual characteristics of the input product design image.
@@
$P_{j}(\widehat{\theta})=\frac{1}{N}\sum_{i\in\mathcal{I}}p_{ij}(\widehat{\theta}),$ (2)
where $\mathcal{I}$ denotes the total consumer set and $N$ is the size of the consumer base. The aggregate popularity measure is later incorporated into the constructed generative models to guide the process of generating new product designs. [^7] Using aggregate popularity labels guides generative models to create designs with higher average popularity. Using individual popularity labels, as demonstrated in the personalization section, enables personalized design generation for specific consumers.
We train two prediction models: one with internal preferences only and another with both internal and external preferences. We train these two prediction models via similar training processes. The difference is that Prediction Model 1 is trained with internal consumer appearances and product designs only, while Prediction Model 2 is trained using both internal and external consumer appearances and product designs. Note that external consumers are used in the training stage of Prediction Model 2 but not in the prediction stage; when using the trained models to calculate the aggregate popularity measures, we average over the choice probabilities of the internal consumers only for both Prediction Models 1 and 2. This is because we are interested in predicting whether the internal consumers of firms prefer the new template designs. The external consumer preference information still enhances the performance of Prediction Model 2 by training the model to recognize popular design features in external designs. The underlying assumption is that external consumers with similar appearances to those of the internal consumers share similar preferences. Therefore, learning the preferences of the external consumers helps us infer internal consumers’ preferences for the new design features from the external designs.
Generative Models
@@
When training the baseline model without consumer preferences, we adopt the standard and well-established model architecture of the DCGAN method (Radford et al.,,2015) to generate new images. This model takes existing internal and external designs as real image inputs and identifies a set of parameters that enables the samples generated from random noises to closely resemble the training data. As the training process is unsupervised and lacks consumer preference information, the generated images are not necessarily preferred by consumers.
When training the enhanced model with internal preferences, we estimate a conditional GAN model and let the internal consumer preferences guide the generator to produce consumer-preferred designs. Specifically, we build on the CcGAN model developed by Ding et al., (2020) which is designed for image generation based on continuous labels and utilizes a novel empirical generator loss and a unique label input method.
@@
$y_{j}=P_{j}(\widehat{\theta}_{1}),$ (3)
where $\widehat{\theta}_{1}$ is the estimated parameter of Prediction Model 1 and $P_{j}(\widehat{\theta}_{1})$ is defined in Equation 2. This popularity measure serves as a continuous class label and is fed into both the generator and the discriminator in the CcGAN to guide the generator when creating new designs. Specifically, building on Ding et al., (2020), we define the hard vicinal discriminator loss (HVDL) as
^$\begin{split}\widehat{\mathcal{L}}(D)=&-\frac{C_{1}}{M^{r}}\sum_{j=1}^{M^{r}}\sum_{i=1}^{M^{r}}\mathbb{E}_{\epsilon^{r}\sim N(0,\sigma^{2})}\Big{[}w_{j}^{r}\cdot\frac{\mathbbm{1}_{\{|y_{j}^{r}+\epsilon^{r}-y_{i}^{r}|\leq\kappa\}}}{N_{y_{j}^{r}+\epsilon^{r},\kappa}^{r}}\mbox{log}(D(\boldsymbol{x}_{i}^{r},y_{j}^{r}+\epsilon^{r}))\Big{]}\\ &-\frac{C_{2}}{M^{g}}\sum_{j=1}^{M^{g}}\sum_{i=1}^{M^{g}}\mathbb{E}_{\epsilon^{g}\sim N(0,\sigma^{2})}\Big{[}w_{j}^{g}\cdot\frac{\mathbbm{1}_{\{|y_{j}^{g}+\epsilon^{g}-y_{i}^{g}|\leq\kappa\}}}{N_{y_{j}^{g}+\epsilon^{g},\kappa}^{r}}\mbox{log}(1-D(\boldsymbol{x}_{i}^{g},y_{j}^{g}+\epsilon^{g}))\Big{]},\end{split}$ (4)
where $C_{1}$ and $C_{2}$ are constants, $\boldsymbol{x}_{i}^{r}$ and $\boldsymbol{x}_{i}^{g}$ stand for real and generated images, respectively, $y_{i}^{r}$ and $y_{i}^{g}$ are the popularity labels of $\boldsymbol{x}_{i}^{r}$ and $\boldsymbol{x}_{i}^{g}$, $y_{j}^{r}$ and $y_{j}^{g}$ are target labels in the datasets, and $\epsilon^{r}$ and $\epsilon^{g}$ are normally distributed random noises. $N_{y_{j}^{r}+\epsilon^{r},k}^{r}$ is the number of the $y_{i}^{r}$ that satisfies $|y_{j}^{r}+\epsilon^{r}-y_{i}^{r}|\leq\kappa$. The label weights {$w_{j}^{r},w_{i}^{g}$} are defined to be proportional to the labels as {$w_{j}^{r}=ky_{j}^{r},w_{j}^{g}=ky_{j}^{g}$}, where $k$ is a hyper-parameter that takes a positive value. The loss function for the generator is defined as
^$\widehat{\mathcal{L}}(G)=-\frac{1}{M^{g}}\sum_{i=1}^{M^{g}}w_{i}^{g}\cdot\mathbb{E}_{\epsilon^{g}\sim N(0,\sigma^{2})}\mbox{log}(D(G(\boldsymbol{z}_{i},y_{i}^{g}+\epsilon^{g}),y_{i}^{g}+\epsilon^{g})),$ (5)
@@
An important modification we made to Ding et al., (2020) is the introduction of label weights, $\{w_{j}^{r},w_{j}^{g},w_{i}^{g}$}, to the loss functions. In Ding’s work, the continuous labels, such as yaw angles of 3-D chair models and human ages, do not imply preference. However, in our context, higher popularity levels indicate greater consumer preference, and we aim to generate designs with higher popularity. To enable CcGAN to recognize the ordinal nature of popularity labels, we include label weights, with higher popularity labels assigned larger weights, indicating their greater importance for the algorithm. As demonstrated in the result sections, incorporating these label weights significantly enhances the performance of the generative model.
In summary, combining Equations 3, 4, and 5, we integrate consumer preferences into the generation process by directly incorporating the popularity measure $P_{j}(\widehat{\theta})$ from the predictor model into the loss functions of both the discriminator and the generator of the CcGAN. This setup allows us to instruct the generator to create new images with a specific popularity level.
The advanced model incorporating both internal and external preferences is trained using a procedure similar to that of the enhanced model. The only difference is that we use the aggregate popularity measures from Prediction Model 2, $P_{j}(\widehat{\theta}_{2})$, as additional input labels to be fed into the generator and the discriminator. As Prediction Model 2 is trained on both internal and external consumer choices, the new CcGAN is able to incorporate new design features from the external sources that are appealing to the internal consumers.
Institutional Setting and Data
Institutional Background
As an empirical application, we apply our proposed framework to a real business setting: a self-operated photography chain in China. The company operates self-operated photo booths in approximately twenty shopping malls in major cities. The photo booths allow consumers to choose photo templates, take photos with the chosen templates, and print their photos. The photo booths are self-operated, so consumers follow the instructions on the photo booth machines. Figure 2 shows the interface of such a photo machine. Four photo categories can be selected: main (single), friends, family, and lovers. We focus on the “main” category in our analysis, which is designed for one consumer to take photos.
@@
Internal Data
The internal dataset obtained from the company contains two parts. The first part includes the original template images that the company offered to customers, without the presence of consumers inside the images. The company offered 83 photo themes in total during the observation period. Each theme contains 1 to 13 templates, yielding a total of 585 templates across all the themes. The second part of the internal dataset contains the complete order histories of 2195 consumers who had taken at least one photo between October 2017 and August 2018. For each consumer, we observe their selected themes and templates and the photos taken with the chosen templates. We also observe the ranking orders of the themes offered on the screen at the times they made their choices.
The internal data patterns suggest that the templates vary significantly in popularity both across themes and within the same theme, potentially due to different design features and heterogeneous consumer preferences. The variation in template popularity, especially the presence of low-popularity templates, suggests that the company needs to improve the attractiveness of its template designs. As another piece of evidence, the internal dataset shows that only 2195 out of the 6038 registered consumers ended up taking photos in the booth after browsing the photo template options. Among the 2195 consumers, fewer than 100 consumers made multiple store visits, and only 15% of the consumers had repeated time purchases. These numbers suggest that the company can benefit from incorporating consumer preferences to generate better designs that are appealing to consumers.
External Data
@@
After gathering the external images with human objects and backgrounds, we preprocess the images before they are used as inputs for the proposed model. Specifically, training the model requires a set of pure templates without human objects and a set of consumer face images, while the raw external images we collect are combinations of template backgrounds and consumer faces, so we need to separate the two components. For each external image, we remove the human object in it and inpaint the remaining part of the image to obtain a complete background image, which is similar to the pure template images in the internal dataset. Additionally, we save the cropped human faces as new images, which are similar to the consumer face images in the internal dataset. Figure 4 presents examples of the internal templates, raw external photos, inpainted external templates, and cropped consumer faces, all from the graffiti theme. As shown in the graphs, the collected external images are similar in style to the internal images and have sufficient variations. The inpainting process works well and can generate high-quality templates and consumer face images.
@@
Applying Model to Practice
@@
We apply the proposed model to generate new product designs for the unmanned photo gallery firm. This section describes the implementation details.
Training the Prediction Model
@@
The feature embeddings of consumer faces and background templates serve as the inputs for the prediction model. Regarding the consumer face features, most consumers took more than one photo, so multiple face images are available for each consumer. We first obtain the embedding of each face image. We then average across the embeddings of the same consumer to obtain a 128-dimensional facial vector $\overrightarrow{X}_{i}$ for that consumer. With respect to the template features, in addition to the 1000-dimensional embeddings of the visual characteristics $\overrightarrow{V}_{j}$, we also include the display ranking of each template on the screen $\overrightarrow{Z}_{j}$, as its display ranking affects its probability of being chosen. Concatenating $\overrightarrow{X}_{i}$, $\overrightarrow{V}_{j}$, and $\overrightarrow{Z}_{j}$ yields a 1129-dimensional feature input for training the prediction model.
[…]
@@
The output variable of the prediction model is constructed from the consumers’ revealed preferences, namely, their template choices. In our setting, consumers first choose a theme among all themes; only one template from each theme is displayed as the cover picture for each theme at this stage. Once consumers click on the chosen theme, they can browse all templates in the chosen theme and choose some of the templates to take photos. We assume that consumers choose among templates they have browsed (i.e., cover-page templates of each theme and all templates within the chosen theme). Each choice observation is a consumer-template pair consisting of one consumer and one template she has browsed. If she chose the template, the observation counts towards a positive sample. If she did not choose, the observation counts towards a negative sample. Consumers make binary choices for each browsed templates and can choose multiple templates since consumers can select an unlimited number of themes and templates within the paid time window. The chosen templates belonging to the chosen theme naturally serve as the positive samples. The unchosen templates belonging to the chosen theme serve as the negative samples because consumers browse them but decide not to take photos with them. For the unchosen themes, the cover-page templates count toward the negative samples because consumers browse them upon choosing the themes and do not choose them. However, the other templates belonging to the unchosen themes do not count as negative samples and are not included in the analysis because consumers do not have a chance to see them. Overall, the output variable $Y_{ij}$ for consumer $i$ and template $j$ is constructed as follows:
@@
$Y_{ij}=\left\{\begin{array}[]{ll}1&\text{if template }j\text{ is chosen by consumer }i\\ 0&\text{if template }j\text{ is not chosen but belongs to a chosen theme}\\ 0&\text{if template }j\text{ is the cover-page template of an unchosen theme }\end{array}\right.$ (7)
Given the definitions of the input and output variables, we estimate Prediction Model 1 using internal consumer preferences only. We focus on photo samples from the category “Main (single),” which includes 931 consumers and 33,804 consumer-template pairs. Each pair consists of one consumer and one template that the consumer has browsed. Out of these 33,804 pairs, 3,629 are positive samples, and 30,175 are negative samples. A positive sample indicates that the consumer chose the template, while a negative sample indicates that the consumer browsed but did not choose the template. Negative samples can include templates from both the unchosen templates of the chosen themes and the cover-page templates of the unchosen themes. It’s important to note that a template can appear in both the positive and negative samples, depending on the consumer it is paired with. In our data, every internal template has been chosen by at least one consumer, meaning each internal template is part of both the negative and positive samples. The company offered 83 photo themes and 585 photo templates during the observation period. Therefore, there are 585 unique internal templates in both the positive and negative samples for training the prediction model.
@@
We retain $40\%$ of the observations as the test set and use the remaining data for training the model. The training and test sets include 20,282 and 13,522 observations, respectively. As positive samples only account for 10.7% of the total sample size, an unbalanced sample problem is encountered. To alleviate this problem, we upsample the minority class to 18,000 to better match the majority class. We concatenate the upsampled minority class and the majority class to form the final training set and train the model using a random forest. The number of trees in the forest is set to 100. The class weights are set to be balanced.
@@
We estimate Prediction Model 2 using both internal and external consumer preferences. Specifically, external UGC data contain photos taken by individual users in front of various backgrounds. Each photo represents a user’s chosen background and thus serves as a positive sample for the prediction model. Taking the graffiti theme as an example. There are 1375 external photos in the graffiti them, each representing a unique consumer and a unique external template. There are 1375 unique consumers and 1375 unique external templates. These 1375 consumer-template pairs serve as positive samples. Combining these with the 33,804 internal observations yields a sample size of 35,179 consumer-template pairs. We use 70% of the observations, or 24,625 pairs, as the training set. Among them, 2,544 are positive samples, and 22,081 are negative samples. To address the unbalanced sample issue, we upsample the minority class to 20,000 to align with the majority class. We estimate Prediction Model 2 using a random forest again, maintaining the same hyperparameters as those employed in Prediction Model 1.
Training the Generative Models and Generating New Designs